PLEASE SHOW THE WORKING!!

MP3 Players CD Plavers 4 4 12 16 20 24

1. The table above shows the production possibilities for two goods, MP3 players and CD players. The opportunity cost of producing the 4th CD player is

16 MP3 players

8 MP3 players

4 MP3 players

2 MP3 players

1 MP3 player

2. After saving money in her piggy bank for 3 years, Beverly decided to deposit $5,000 of the money in the Millertown bank. If the bank were fully “loaned out’’ and the required reserve ratio were 20%, then the maximum change in the money supply due to this deposit would be

$ 25,000

$ 20,000

$ 5,000

$ 4, 000

3. If you were a classical economist, which of the following is more consistent with your beliefs concerning an increase in the money supply?

An increase in the money supply will cause a decrease in price level and an increase in nominal wages.

An increase in the money supply will cause a decrease in price level and nominal wages.

An increase in the money supply will cause an increase in price level and nominal wages.

An increase in the money supply will cause an increase in price level and a decrease in nominal wages.

An increase in the money supply will not impact the economy
Answer #1

1. Ans: 4 MP3 player

Explanation:

When 4th CD player is produced, MP3 player production is reduced to 8 units from 12 units. It means the opportunity cost of producing the 4th CD player is 4 (i.e., 12 - 8) MP3 player.

2. Ans: $25,000

Explanation:

Money multiplier = 1 / required reserve ratio = 1 / 0.2 = 5

Maximum change in the money supply = $5,000 * 5 = $25,000

3. Ans: An increase in the money supply will cause an increase in price level and nominal wages.